# Representations of a quantum-deformed Lorentz algebra, Clebsch-Gordan map, and Fenchel-Nielsen representation of quantum complex flat connections at level-$\bm{k}$

### v1

@@ 4.2 S-cycle trace operator
$e^{L_{3}}=\exp\left[\frac{2\pi i}{k}\left(-ib\mu_{3}-m_{3}\right)\right],\qquad e^{L_{4}}=\exp\left[\frac{2\pi i}{k}\left(-ib\mu_{4}-m_{4}\right)\right]$ (4.39)
where $\mu_{3},\mu_{4}\in\mathbb{R}$ and $m_{3},m_{4}\in\mathbb{Z}$. For this paper, the purpose of this assumption is to simplify some analysis in the following. However, when we consider the quantization of $\mathrm{SL}(2,\mathbb{C})$ Chern-Simons theory on 3-manifold whose boundary is a closed 2-surface containing the 4-holed sphere as a part of the geodesic boundary (see e.g. [22, 18]), $\{L_{a}\}_{a=1}^{4}$ are part of the Darboux coordinate of the phase space, so $\{e^{L_{a}}\}_{a=1}^{4}$ should be quantized in the same way as $e^{\bm{Y}}$.
We apply the unitary shift operator
$\mathcal{S}_{34}=\exp\left[-\frac{2\pi i}{k}\left((\mu_{3}+\mu_{4})\bm{\nu}-(m_{3}+m_{4})\bm{n}\right)\right]$ (4.40)
@@
$\mathcal{S}_{34}e^{-{\bm{Y}}+L_{3}+L_{4}}\mathcal{S}_{34}^{-1}=e^{-{\bm{Y}}}$ (4.41)
similar to (3.24). We obtain
$\mathcal{S}_{34}\mathcal{F}^{-1}\bm{L}_{s}^{\prime}\mathcal{F}\mathcal{S}_{34}^{-1}=e^{-2{\bm{X}}+L_{3}+L_{4}}+e^{2{\bm{X}}-L_{3}-L_{4}}+e^{-{\bm{Y}}},$ (4.42)
which relates to $\mathcal{Q}_{1}^{\prime\prime}$ by identifying $L_{3}+L_{4}$ here to $L_{2}$ in $\mathcal{Q}_{1}^{\prime\prime}$. The relation between $\tilde{\bm{L}}_{s}^{\prime}$ and $\tilde{\mathcal{Q}}_{1}^{\prime\prime}$ is obtained analogously.
If we further assume $m_{3}+m_{4}\in 2\mathbb{Z}$, we can use another unitary shift operator
$\mathcal{D}_{34}=\exp\left[-\frac{2\pi i}{k}\left(\frac{\mu_{3}+\mu_{4}}{2}\bm{\mu}-\frac{m_{3}+m_{4}}{2}\bm{m}\right)\right]$ (4.43)
to give
$\mathcal{D}_{34}^{-1}e^{2\bm{X}-L_{3}-L_{4}}\mathcal{D}_{34}=e^{2\bm{X}},\qquad\mathcal{D}_{34}^{-1}e^{-2\bm{X}+L_{3}+L_{4}}\mathcal{D}_{34}=e^{-2\bm{X}}.$ (4.44)
As a result, the sequence of unitary transformation gives a relatively simple expression of S-cycle trace operator
$\bm{L}_{s}^{\prime\prime}=\mathcal{D}_{34}^{-1}\mathcal{S}_{34}\mathcal{F}^{-1}\bm{L}_{s}^{\prime}\mathcal{F}\mathcal{S}_{34}^{-1}\mathcal{D}_{34}=e^{-2{\bm{X}}}+e^{2{\bm{X}}}+e^{-{\bm{Y}}}$ (4.45)
$\tilde{\bm{L}}_{s}^{\prime\prime}=\mathcal{D}_{34}^{-1}\mathcal{S}_{34}\mathcal{F}^{-1}\tilde{\bm{L}}_{s}^{\prime}\mathcal{F}\mathcal{S}_{34}^{-1}\mathcal{D}_{34}=e^{-2{\tilde{\bm{X}}}}+e^{2\tilde{\bm{X}}}+e^{-\tilde{\bm{Y}}}.$ (4.46)
In summary, we have made a sequence of unitary transformations
$\mathfrak{U}_{s}=\mathcal{D}_{34}^{-1}\mathcal{S}_{34}\mathcal{F}^{-1}\bm{U}_{s}$ (4.47)
so that
$\bm{L}_{s}^{\prime\prime}=\mathfrak{U}_{s}\bm{L}_{s}\mathfrak{U}_{s}^{-1},\qquad\tilde{\bm{L}}_{s}^{\prime\prime}=\mathfrak{U}_{s}\tilde{\bm{L}}_{s}\mathfrak{U}_{s}^{-1}.$ (4.48)
@@ 4.3 T-cycle trace operator
$\bm{L}_{t}$ $=$ $e^{-2\bm{Y}_{t}}+e^{2\bm{Y}_{t}}+e^{-L_{1}+L_{2}-L_{3}-L_{4}-\bm{X}+2\bm{Y}_{t}}+e^{-L_{1}-L_{4}-\bm{X}}+e^{L_{2}-L_{3}-\bm{X}}+e^{-\bm{X}-2\bm{Y}_{t}}$ (4.49)
$=$ $e^{-2\bm{Y}_{t}}+e^{2\bm{Y}_{t}}+e^{L_{2}+L_{3}}{\bf q}e^{-2\bm{Y}}\left(1+e^{-2L_{3}}e^{2\bm{Y}}{\bf q}^{-1}\right)\left(1+e^{-\left(L_{1}+L_{2}+L_{3}+L_{4}\right)}e^{2\bm{Y}}{\bf q}^{-1}\right)e^{-\bm{X}}$
where $2\bm{Y}_{t}=2\bm{Y}-L_{2}-L_{3}$. We obtain the tilded operator by $\tilde{\bm{L}}_{t}=\bm{L}_{t}^{\dagger}$. We apply the unitary transformation
$\bm{U}^{\prime}_{t}\bm{L}_{t}\bm{U}^{\prime-1}_{t}$ $=$ $e^{-2\bm{Y}+L_{2}+L_{3}}+e^{2\bm{Y}-L_{2}-L_{3}}+e^{-2\bm{Y}-\bm{X}+L_{2}+L_{3}},$ (4.50)
@@
$\bm{U}^{\prime}_{t}$ $=$ $\varphi\left(e^{-\left(L_{1}+L_{2}+L_{3}+L_{4}\right)}y^{2},e^{-\left(\tilde{L}_{1}+\tilde{L}_{2}+\tilde{L}_{3}+\tilde{L}_{4}\right)}\tilde{y}^{2}\right)^{-1}\varphi\left(e^{-2L_{3}}y^{2},e^{-2\tilde{L}_{3}}\tilde{y}^{2}\right)^{-1}.$ (4.51)
The further simplification can be made by the Weil transformation representing the following symplectic transformation
$\left(\begin{array}[]{c}Y\\ X\end{array}\right)\to\begin{pmatrix}-1&0\\ -2&-1\end{pmatrix}\left(\begin{array}[]{c}Y\\ X\end{array}\right)=\begin{pmatrix}1&0\\ 2&1\end{pmatrix}\begin{pmatrix}-1&0\\ 0&-1\end{pmatrix}\left(\begin{array}[]{c}Y\\ X\end{array}\right)$ (4.64)
@@
$\bm{L}_{t}^{\prime}=\bm{U}_{t}\bm{L}_{t}\bm{U}_{t}^{-1}=e^{-2\bm{Y}-L_{2}-L_{3}}+e^{2\bm{Y}+L_{2}+L_{3}}+e^{\bm{X}+L_{2}+L_{3}},\qquad\bm{U}_{t}=\bm{T}^{-1}\bm{V}\bm{U}^{\prime}_{t}.$ (4.65)
Similar to $\bm{L}_{s}$, we further make the Fourier transform:
$\mathcal{F}^{-1}\bm{L}_{t}^{\prime}\mathcal{F}=e^{-2{\bm{X}}-L_{2}-L_{3}}+e^{2{\bm{X}}+L_{2}+L_{3}}+e^{-{\bm{Y}}+L_{2}+L_{3}}.$ (4.66)
If both $e^{L_{2}},e^{L_{3}}$ can be parametrized by
$e^{L_{2}}=\exp\left[\frac{2\pi i}{k}\left(-ib\mu_{2}-m_{2}\right)\right],\qquad e^{L_{3}}=\exp\left[\frac{2\pi i}{k}\left(-ib\mu_{3}-m_{3}\right)\right]$ (4.67)
where $\mu_{2},\mu_{3}\in\mathbb{R}$ and $m_{2},m_{3}\in\mathbb{Z}$. Apply the shift operator
$\mathcal{S}_{23}=\exp\left[-\frac{2\pi i}{k}\left((\mu_{2}+\mu_{3})\bm{\nu}-(m_{2}+m_{3})\bm{n}\right)\right]$ (4.68)
@@
$\mathcal{S}_{23}\mathcal{F}^{-1}\bm{L}_{t}^{\prime}\mathcal{F}\mathcal{S}_{23}^{-1}=e^{-2{\bm{X}}-L_{2}-L_{3}}+e^{2{\bm{X}}+L_{2}+L_{3}}+e^{-{\bm{Y}}},$ (4.69)
which relates to $\mathcal{Q}_{1}^{\prime\prime}$ by identifying $-L_{2}-L_{3}$ here to $L_{2}$ in $\mathcal{Q}_{1}^{\prime\prime}$.
If we assume $m_{2}+m_{3}\in 2\mathbb{Z}$, we can define another shift operator
$\mathcal{D}_{23}=\exp\left[-\frac{2\pi i}{k}\left(\frac{\mu_{2}+\mu_{3}}{2}\bm{\mu}-\frac{m_{2}+m_{3}}{2}\bm{m}\right)\right]$ (4.70)
which gives
$\mathcal{D}_{23}e^{2\bm{X}+L_{2}+L_{3}}\mathcal{D}_{23}^{-1}=e^{2\bm{X}},\qquad\mathcal{D}_{23}e^{-2\bm{X}-L_{2}-L_{3}}\mathcal{D}_{23}^{-1}=e^{-2\bm{X}}.$ (4.71)
As a result, the sequence of unitary transformation gives same expression as (4.45) for the S-cycle trace operator
$\bm{L}_{t}^{\prime\prime}=\mathcal{D}_{23}\mathcal{S}_{23}\mathcal{F}^{-1}\bm{L}_{t}^{\prime}\mathcal{F}\mathcal{S}_{23}^{-1}\mathcal{D}_{23}=e^{-2{\bm{X}}}+e^{2{\bm{X}}}+e^{-{\bm{Y}}}=\bm{L}_{s}^{\prime\prime}$ (4.72)
$\tilde{\bm{L}}_{t}^{\prime\prime}=\mathcal{D}_{23}^{-1}\mathcal{S}_{23}\mathcal{F}^{-1}\tilde{\bm{L}}_{t}^{\prime}\mathcal{F}\mathcal{S}_{23}^{-1}\mathcal{D}_{23}^{-1}=e^{-2{\tilde{\bm{X}}}}+e^{2\tilde{\bm{X}}}+e^{-\tilde{\bm{Y}}}=\tilde{\bm{L}}_{s}^{\prime\prime}$ (4.73)
We denote by
$\mathfrak{U}_{t}=\mathcal{D}_{23}\mathcal{S}_{23}\mathcal{F}^{-1}\bm{U}_{t},$ (4.74)
so that
$\bm{L}_{t}^{\prime\prime}=\mathfrak{U}_{t}\bm{L}_{t}\mathfrak{U}_{t}^{-1},\qquad\tilde{\bm{L}}_{t}^{\prime\prime}=\mathfrak{U}_{t}\tilde{\bm{L}}_{t}\mathfrak{U}_{t}^{-1}.$ (4.75)
Then $\bm{L}_{t}^{\prime\prime}=\bm{L}_{s}^{\prime\prime}$, $\tilde{\bm{L}}_{t}^{\prime\prime}=\tilde{\bm{L}}_{s}^{\prime\prime}$ implies that there is a unitary transformation relating $(\bm{L}_{t},\tilde{\bm{L}}_{t})$ and $(\bm{L}_{s},\tilde{\bm{L}}_{s})$.
@@
Finally, an additional Fourier transform gives the final expression of the trace operator that is studied extensively in the next section
$\bm{L}=\mathcal{F}\bm{L}_{s}^{\prime\prime}\mathcal{F}^{-1}=e^{-2{\bm{Y}}}+e^{2{\bm{Y}}}+e^{{\bm{X}}},$ (4.77)
$\tilde{\bm{L}}=\mathcal{F}\tilde{\bm{L}}_{s}^{\prime\prime}\mathcal{F}^{-1}=e^{-2\tilde{\bm{Y}}}+e^{2\tilde{\bm{Y}}}+e^{\tilde{\bm{X}}}.$ (4.78)

### v2

@@ 4.3 S-cycle trace operator
$e^{L_{3}}=\exp\left[\frac{2\pi i}{N}\left(-ib\mu_{3}-m_{3}\right)\right],\qquad e^{L_{4}}=\exp\left[\frac{2\pi i}{N}\left(-ib\mu_{4}-m_{4}\right)\right]$ (4.43)
where $\mu_{3},\mu_{4}\in\mathbb{R}$ and $m_{3},m_{4}\in\mathbb{Z}/N\mathbb{Z}$. For this paper, the purpose of this assumption is to further simplify $\bm{L}_{s}^{\prime}$ and $\tilde{\bm{L}}_{s}^{\prime}$ by unitary transformations and relate them to $\bm{L},\tilde{\bm{L}}$ derived at the end of Section 3.2. However, when we consider the quantization of $\mathrm{SL}(2,\mathbb{C})$ Chern-Simons theory on 3-manifold whose boundary is a closed 2-surface containing the 4-holed sphere as a part of the geodesic boundary (see e.g. [26, 19]), $\{L_{a}\}_{a=1}^{4}$ are part of the Darboux coordinate of the phase space, so $\{e^{L_{a}}\}_{a=1}^{4}$ should be quantized in the same way as $e^{\bm{Y}}$.
We apply the unitary shift operators
$\mathcal{S}_{34}f(\mu,m)=f\left(\mu+\mu_{3}+\mu_{4},m+m_{3}+m_{4}\right),\qquad\mathcal{D}_{34}=e^{\frac{2\pi i}{N}\left[({\mu_{3}+\mu_{4}})\bm{\mu}-({m_{3}+m_{4}})\bm{m}\right]}$ (4.44)
@@
$\mathcal{S}_{34}e^{-{\bm{Y}}+L_{3}+L_{4}}\mathcal{S}_{34}^{-1}=e^{-{\bm{Y}}},\qquad\mathcal{S}_{34}e^{{\bm{U}}}\mathcal{S}_{34}^{-1}=e^{{\bm{U}}}$ (4.45)
$\mathcal{D}_{34}^{-1}e^{\bm{U}+L_{3}+L_{4}}\mathcal{D}_{34}=e^{\bm{U}},\qquad\mathcal{D}_{34}^{-1}e^{\bm{Y}}\mathcal{D}_{34}=e^{\bm{Y}}$ (4.46)
Therefore
$\mathcal{D}_{34}^{-1}\mathcal{S}_{34}\bm{L}_{s}^{\prime}\mathcal{S}_{34}^{-1}\mathcal{D}_{34}$ $=$ $e^{-{\bm{Y}}}+e^{{\bm{Y}}}+e^{{\bm{U}}}=\bm{L},$ (4.47)
$\mathcal{D}_{34}^{-1}\mathcal{S}_{34}\tilde{\bm{L}}_{s}^{\prime}\mathcal{S}_{34}^{-1}\mathcal{D}_{34}$ $=$ $e^{-{\tilde{\bm{Y}}}}+e^{\tilde{\bm{Y}}}+e^{\tilde{\bm{U}}}=\tilde{\bm{L}}.$ (4.48)
This shows that the sequence of unitary transformations on $\mathcal{H}$
$\mathfrak{U}_{s}=\mathcal{D}_{34}^{-1}\mathcal{S}_{34}\bm{U}_{s}$ (4.49)
relates the S-cycle trace operators $\bm{L}_{s},\tilde{\bm{L}}_{s}$ to $\bm{L},\tilde{\bm{L}}$ derived in Section 3.2 from the Clebsch-Gordan decomposition of $\mathscr{U}_{\mathbf{q}}(sl_{2})\otimes\mathscr{U}_{\widetilde{\mathbf{q}}}(sl_{2})$ unitary representations:
$\bm{L}=\mathfrak{U}_{s}\bm{L}_{s}\mathfrak{U}_{s}^{-1},\qquad\tilde{\bm{L}}=\mathfrak{U}_{s}\tilde{\bm{L}}_{s}\mathfrak{U}_{s}^{-1}.$ (4.50)
@@ 4.4 T-cycle trace operator
$\bm{L}_{t}$ $=$ $e^{-\bm{Y}_{t}}+e^{\bm{Y}_{t}}-e^{L_{2}+L_{3}}{\bf q}e^{-\bm{Y}}\left(1+{\bf q}^{-1}e^{-2L_{3}+\bm{Y}}\right)\left(1+{\bf q}^{-1}e^{-L_{1}-L_{2}-L_{3}-L_{4}+\bm{Y}}\right)e^{-\bm{U}}$ (4.51)
where $\bm{Y}_{t}=\bm{Y}-L_{2}-L_{3}$. We obtain the tilded operator by $\tilde{\bm{L}}_{t}=\bm{L}_{t}^{\dagger}$. We apply the unitary transformation:
$\bm{U}^{\prime}_{t}\bm{L}_{t}\bm{U}^{\prime-1}_{t}$ $=$ $e^{-\bm{Y}+L_{2}+L_{3}}+e^{\bm{Y}-L_{2}-L_{3}}-e^{-\bm{Y}-\bm{U}+L_{2}+L_{3}},$ (4.52)
@@
$\bm{U}^{\prime}_{t}$ $=$ $\varphi\left(e^{-\left(L_{1}+L_{2}+L_{3}+L_{4}\right)}\bm{y},e^{-\left(\tilde{L}_{1}+\tilde{L}_{2}+\tilde{L}_{3}+\tilde{L}_{4}\right)}\tilde{\bm{y}}\right)^{-1}\varphi\left(e^{-2L_{3}}\bm{y},e^{-2\tilde{L}_{3}}\tilde{\bm{y}}\right)^{-1}.$ (4.53)
The further unitary transformation can be made by the Weil transformation representing the following symplectic transformation
$\left(\begin{array}[]{c}Y\\ U\end{array}\right)\to\begin{pmatrix}-1&0\\ -1&-1\end{pmatrix}\left(\begin{array}[]{c}Y\\ U\end{array}\right)=\begin{pmatrix}1&0\\ 1&1\end{pmatrix}\begin{pmatrix}-1&0\\ 0&-1\end{pmatrix}\left(\begin{array}[]{c}Y\\ U\end{array}\right).$ (4.60)
@@
$\bm{L}_{t}^{\prime}=\bm{U}_{t}\bm{L}_{t}\bm{U}_{t}^{-1}=e^{-\bm{Y}-L_{2}-L_{3}}+e^{\bm{Y}+L_{2}+L_{3}}+e^{\bm{U}+L_{2}+L_{3}},\qquad\bm{U}_{t}=\bm{T}^{-1}\bm{V}\bm{U}^{\prime}_{t}.$ (4.61)
Similar to $\bm{L}_{s}$, we parametrize $e^{L_{2}},e^{L_{3}}$ by
$e^{L_{2}}=\exp\left[\frac{2\pi i}{N}\left(-ib\mu_{2}-m_{2}\right)\right],\qquad e^{L_{3}}=\exp\left[\frac{2\pi i}{N}\left(-ib\mu_{3}-m_{3}\right)\right]$ (4.62)
where $\mu_{2},\mu_{3}\in\mathbb{R}$ and $m_{2},m_{3}\in\mathbb{Z}/N\mathbb{Z}$. We apply the unitary shift operators
$\mathcal{S}_{23}f(\mu,m)=f\left(\mu+\mu_{2}+\mu_{3},m+m_{2}+m_{3}\right),\qquad\mathcal{D}_{23}=e^{\frac{2\pi i}{N}\left[({\mu_{2}+\mu_{3}})\bm{\mu}-({m_{2}+m_{3}})\bm{m}\right]},$ (4.63)
@@
$\mathcal{S}_{23}^{-1}e^{{\bm{Y}}+L_{2}+L_{3}}\mathcal{S}_{23}$ $=$ $e^{{\bm{Y}}},\qquad\mathcal{S}_{23}^{-1}e^{{\bm{U}}}\mathcal{S}_{23}=e^{{\bm{U}}}$ (4.64)
$\mathcal{D}_{23}^{-1}e^{\bm{U}+L_{2}+L_{3}}\mathcal{D}_{23}$ $=$ $e^{\bm{U}},\qquad\mathcal{D}_{23}^{-1}e^{\bm{Y}}\mathcal{D}_{23}=e^{\bm{Y}}$ (4.65)
As a result, the sequence of unitary transformations transforms the T-cycle trace operators to $\bm{L},\tilde{\bm{L}}$
$\mathcal{D}_{23}^{-1}\mathcal{S}_{23}^{-1}\bm{L}_{t}^{\prime}\mathcal{S}_{23}\mathcal{D}_{23}$ $=$ $e^{-{\bm{Y}}}+e^{{\bm{Y}}}+e^{{\bm{U}}}=\bm{L},$ (4.66)
$\mathcal{D}_{23}^{-1}\mathcal{S}_{23}\tilde{\bm{L}}_{t}^{\prime}\mathcal{S}_{23}\mathcal{D}_{23}$ $=$ $e^{-{\tilde{\bm{Y}}}}+e^{\tilde{\bm{Y}}}+e^{\tilde{\bm{U}}}=\tilde{\bm{L}}.$ (4.67)
We denote the unitary operator $\mathfrak{U}_{t}$ by
$\mathfrak{U}_{t}=\mathcal{D}_{23}^{-1}\mathcal{S}_{23}^{-1}\bm{U}_{t},$ (4.68)
so the above analysis shows that
$\bm{L}=\mathfrak{U}_{t}\bm{L}_{t}\mathfrak{U}_{t}^{-1},\qquad\tilde{\bm{L}}=\mathfrak{U}_{t}\tilde{\bm{L}}_{t}\mathfrak{U}_{t}^{-1},$ (4.69)
and it also implies that there is a unitary transformation relating $(\bm{L}_{t},\tilde{\bm{L}}_{t})$ and $(\bm{L}_{s},\tilde{\bm{L}}_{s})$.